OpenStudy (anonymous):

Write an equation of a line parallel to y = 2 over 3x −5 in slope-intercept form that passes through the point (−6, −1).

OpenStudy (imstuck):

This is weird. In order to find the slope of the line you gave I graphed it on my calculator. The slope is 5/8. I couldn't figure out any other way to get that particular equation into a y = mx + b format. Anyway, using the point-slope format, we get\[y+1=\frac{ 5 }{ 8 }(x+6)\]Distribute the 5/8 into the parenthesis and get\[y+1=\frac{ 5 }{ 8 }x+\frac{ 30 }{ 8 }\]Now subtract the 1 (in the form 8/8) from both sides.\[y=\frac{ 5 }{ 8 }x+\frac{ 30 }{ 8 }-\frac{ 8 }{ 8 }\]Simplifying gives you\[y=\frac{ 5 }{ 8 }x+\frac{ 22 }{ 8 }\]which reduces to \[y=\frac{ 5 }{ 8 }x+\frac{ 11 }{ 4 }\]That's the best I can do with that crazy first equation!!

OpenStudy (anonymous):

Thank You very much your a lifesaver